1. The area of a rectangular card is 128cm^2. The length of the card is twice its width. What are the dimensions of the card?

Answer by mananth(16949)

let width of card be w
length = 2w
..
Area = L*W
w*2w= 128
2w^2=128
/2
w^2= 64
take square root
w= 8 cm..... thw width
length = 2w = 16 cm
...
m.ananth@hotmail.ca